OpenStudy (anonymous):

Lola has 7 T-shirts. She packed 2 of the T-shirts to bring on vacation. What percent of her T-shirts did Lola bring on vacation? Round to the nearest whole percent.

OpenStudy (nirmalnema):

2/7 *100=200/7 =28.5%